OpenStudy (helpmeplease!!!):

what is 2 out of 198 in percent form?

OpenStudy (stamp):

\[2/198=x\%\]\[2/198=x/100\]\[x=2/198*100\]

OpenStudy (helpmeplease!!!):

I dont get it :/

OpenStudy (helpmeplease!!!):

is it 1%?

OpenStudy (stamp):

@HelpMePlease!!! ~ 1%. 200/198 = 1 1/99 = 1.01010101010101010101...

OpenStudy (helpmeplease!!!):

Thanks @stamp

Directrix (directrix):

@HelpMePlease!!! You can go over to Google.com and enter (2/198) times 100 in the search box. The Google calculator will pop up with 1.01 to which you will add %. So, 1.01% is the approximate answer. Hope that helps a bit.

OpenStudy (helpmeplease!!!):

Thanks @Directrix

Directrix (directrix):

We are all happy to help.

OpenStudy (helpmeplease!!!):

:)